A car starts from rest and accelerates uniformly at a rate of 2 m.s-2 for 20 s. It then maintains a constant velocity for 10 s. The brakes are then applied, and the car is uniformly decelerating and comes to rest in 5 s. What is the maximum velocity and the deceleration in the last 5 seconds?

Question 1 options:


a) 

20 m/s and 4 m/s2.


b) 

10 m/s and 2 m/s2.


c) 

No option is correct.


d) 

30 m/s and 6m/s2.


e) 

40 m/s and 8 m/s2.



1
Expert's answer
2021-07-15T10:51:46-0400

The maximum velocity

"v_2=v_1+at_1=0+2*20=40\:\rm m/s"

The deceleration

"a_3=\frac{v_3-v_2}{\Delta t_3}=\frac{0-40}{5}=-8\:\rm m/s^2"

Answer: e)